# Characterisation of the big Ramsey degrees of the generic partial order

### v4

@@ 8. The lower bound
We briefly discuss the second part of Theorem 1.7, namely the existence of a big Ramsey structure and the connections to topological dynamics. All terminology pertaining to topological dynamics which is not defined here can be found in the introduction of [43]. As we mentioned above, the monoid $\mathrm{Emb}(\mathbf{P},\mathbf{P})$ acts on the right on the set $r^{\mathrm{Emb}(\mathbf{Q},\mathbf{P})}$. It is natural to equip both $\mathrm{Emb}(\mathbf{P},\mathbf{P})$ and $r^{\mathrm{Emb}(\mathbf{Q},\mathbf{P})}$ with topologies as follows.
@@
A completion flow for $G$ is a $G$-flow $X$ containing a completion point, some $x\in X$ such that for any $\eta\in\widehat{G}$, the point $x\cdot\eta\in X$ has dense $G$-orbit, equivalently dense $\widehat{G}$-orbit. One of the main theorems from [43] is that if $G=\mathrm{Aut}(\mathbf{K})$ for $\mathbf{K}$ a Fraïssé structure which admits a big Ramsey structure, then $G$ admits a completion flow $X$ which is universal, i.e. which admits a $G$-map onto any other completion flow $Y$. For $G$ of this form, we further have that the universal completion flow is both metrizable and unique up to isomorphism.
@@
Let $f\in\mathrm{Emb}((\Sigma^{*},\preceq),(\Sigma^{*},\preceq))$. Let $g\in\mathrm{Shape}\allowbreak(\Sigma^{*},\allowbreak\Sigma^{*})$ and sequence $(N_{i})_{i\in\omega}$ be given by Lemma 8.3. Put $f^{\prime}=f\circ g\circ d$. Then for every poset-diary $S$ and every $\ell<\sup_{u\in S}|u|$, we have that $(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)\cong(\overline{f^{\prime}[S]}_{N_{3\ell}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$.
@@
$(\overline{S}_{0},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)=(\overline{f^{\prime}[S]}_{N_{0}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq).$
Now assume that $(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ is isomorphic to $(\overline{f^{\prime}[S]}_{N_{3\ell}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$. We define the function $\mu$ on $\overline{S}_{\ell+1}$ via
@@
We claim that $\mu\colon(\overline{S}_{\ell+1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)\to(\overline{f^{\prime}[S]}_{N_{3\ell+3}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ is an isomorphism:
@@
Fix $f\in\mathrm{Emb}(\mathbf{P},\mathbf{P})$. Let $\psi\colon\mathbf{P}\to{(\Sigma^{*},\preceq)}$ be obtained by Theorem 5.2. Let $\eta\colon{(\Sigma^{*},\preceq)}\to\mathbf{P}$ be an embedding (which exists since $\mathbf{P}$ is universal). Now $\psi\circ f\circ\eta$ is an embedding ${(\Sigma^{*},\preceq)}\to{(\Sigma^{*},\preceq)}$. Let $g\colon{(\Sigma^{*},\preceq)}\to{(\Sigma^{*},\preceq)}$ and $(N_{i})_{i\in\omega}$ be obtained by the application of Lemma 8.3 on $\psi\circ f\circ\eta$. Put $f^{\prime}=\psi\circ f\circ\eta\circ g\circ d$. We claim that for every poset-diary $S$, we have that $\tau(f^{\prime}[S])=S$. By Lemma 8.4, we know that for every $\ell<\sup_{u\in S}|u|$, we have that $(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)\cong(\overline{f^{\prime}[S]}_{N_{3\ell}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$. By Proposition 4.1, for every $0<\ell<\sup_{u\in S}|u|$, there is exactly one difference between $(\overline{S}_{\ell-1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ and $(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$. Consequently there is only one interesting level $\ell^{\prime}$ of $\overline{f^{\prime}[S]}$ between $N_{3\ell}$ and $N_{3\ell+3}$, and $(\overline{f^{\prime}[S]}_{N_{\ell^{\prime}}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)\cong(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ while $(\overline{f^{\prime}[S]}_{N_{\ell^{\prime}+1}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)\cong(\overline{S}_{\ell+1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$. We further note that by the construction of $g$, the map $\tau_{f^{\prime}[S]}\circ f^{\prime}$ must be the identity on $S$. We can thus put $\varphi=\eta\circ g\circ d$. ∎
@@
Given a finite poset $\mathbf{A}$, the fact that the big Ramsey degree of $\mathbf{A}$ is exactly $|T^{\mathrm{lab}}(\mathbf{A})|=|T(\mathbf{A})|\cdot|\mathrm{Aut}(\mathbf{A})|$ follows from Corollary 7.8 and Theorem 8.1. To conclude that $\mathbf{P}$ admits a big Ramsey structure, we observe that the colourings $\chi_{\mathbf{A}}$ as $\mathbf{A}$ ranges over all finite posets satisfy the hypotheses of Theorem 7.1 from [43]. Theorem 1.6 from [43] then shows that $\mathrm{Aut}(\mathbf{P})$ admits a metrizable universal completion flow. ∎
@@ 9.1. Comparison to big Ramsey degrees of the order of rationals
Definition 9.1 (Devlin embedding type [16], see also [41], Definition 6.9).
A set $S\subseteq\{\mathrm{L},\mathrm{R}\}^{*}$ is called a Devlin embedding type, if no member of $S$ extends any other and precisely one of the following four conditions is satisfied for every level $\ell$ with $0\leq\ell<\sup_{w\in S}|w|$:
@@
    $\begin{split}\qquad\overline{S}_{\ell+1}&=\{z\in\overline{S}_{\ell}:z\mathbin{<_{\mathrm{lex}}}w\}^{\frown}\mathrm{L}\\ &\qquad\cup\{w^{\frown}\mathrm{L},w^{\frown}\mathrm{R}\}\\ &\qquad\cup\{z\in\overline{S}_{\ell}:w\mathbin{<_{\mathrm{lex}}}z\}^{\frown}\mathrm{R}.\end{split}$
@@
Theorem 9.3 (Devlin [16], see also [41]).
@@
Any infinite Devlin tree whose leaves code the rational order is a big Ramsey structure for the rationals [43].
@@ 9.2. The triangle-free graph
We outline how the techniques introduced in this paper can yield a particularly compact characterisation of the big Ramsey degrees of the generic triangle-free graph. This is a special case of the main result of [3] (see Example 6.2.10). However, we now give a more compact description which is similar to Definition 1.5.
We fix an alphabet $\Sigma=\{0,1\}$ and denote by $\Sigma^{*}$ the set of all finite words in the alphabet $\Sigma$, by $\mathbin{\leq_{\mathrm{lex}}}$ their lexicographic order, and by $|w|$ the length of the word $w$ (whose letters are indexed by natural numbers starting at 00). We will denote the empty word by $\emptyset$.
We consider the following triangle-free graph $\mathbf{G}^{\triangle}$ introduced in [20].
@@
    $\overline{S}_{i+1}=\{z\in S_{i}\setminus\{w\}:z\perp w\}^{\frown}0\cup\{z\in S_{i}\setminus\{w\}:z\not\perp w\}^{\frown}1.$
@@
    $\overline{S}_{i+1}=(S_{i}\setminus\{0^{i}\})^{\frown}0\cup\{{0^{i}}^{\frown}1\}.$
@@
Tringle-free diares use 00 to represent non-edges and $1$ to represent edges using the “passing number” representation which has been known for the Rado graph [41, 27]. It is again interesting to understand triangle-free diaries in the context of their level structures. While for the Rado graph the level structure is trivial and holds no information except for the lexicographic order, in the triangle-free graph it distinguishes two types of vertices. If vertex has no neighbour defined yet (it is a word $0\cdots 0$), it may split into vertices connected by an edge. There is at most one such vertex in a level strcture. On the other hand, if vertex has a neighbour already defined, splitting it to two vertices connected by an edge would close a triangle. Consequently, it is an interesting event when the first neighbour of a vertex is introduced. Moreover, a given pair of vertices may or may not have a common neighbour. If common neighbour exists, there will be no edges between their respective successors. This is recorded by relation $\perp$ and the corresponding interesting event.
Given a triangle-free graph $\mathbf{H}$, we let $T^{\triangle}(\mathbf{H})$ be the set of all triangle-free-types $S$ such that the structure induced by $\mathbf{G}^{\triangle}$ on $L(S)$ is isomorphic to $\mathbf{H}$. We can now recover the following result of [3].
Theorem 9.8 ([3]).
@@ 10. Concluding remarks
The techniques presented in this paper are quite general and can be applied to a relatively wide class of structures in finite relational languages with relations of arity at most 2, defined by special triangle conditions isolated in [5]. Important examples of such structures are metric spaces with finitely many distances, or metric spaces associated to primitive metrically homogeneous graphs of finite diameter from Cherlin’s catalogue of metrically homogeneous graphs [12, 1].
However, forbidding a configuration on 4 or more vertices (such as in the universal homogeneous $K_{4}$-free graph) will run into a limitation of regular trees used here. This can be overcome in two ways. Either a Ramsey theorem for coding trees can be proved [17, 19, 44, 3] or a theorem for trees with finite but unbounded branching can be used [4] to represent a more regular tree constructed from a kind of amalgamation of all coding trees corresponding to all possible enumerations of the structure, as used in this paper.
Structures with relations of arity 3 and more require trees with unbounded branching together with more involved trees of weak types. Upper bounds for structures with very simple restrictions can be obtained using the product form of the Milliken tree theorem [7, 10], while more general results can be obtained using trees with a successor operation [4], see also the announcement in [2]. Even the characterisation of big Ramsey degrees of 3-uniform hypergraphs shows new and somewhat surprising phenomena, and is a topic of a follow-up paper.

### v5

@@ 8. The lower bound
We briefly discuss the second part of Theorem 1.7, namely the existence of a big Ramsey structure and the connections to topological dynamics. All terminology pertaining to topological dynamics which is not defined here can be found in the introduction of [Zuc19]. As we mentioned above, the monoid $\mathrm{Emb}(\mathbf{P},\mathbf{P})$ acts on the right on the set $r^{\mathrm{Emb}(\mathbf{Q},\mathbf{P})}$. It is natural to equip both $\mathrm{Emb}(\mathbf{P},\mathbf{P})$ and $r^{\mathrm{Emb}(\mathbf{Q},\mathbf{P})}$ with topologies as follows.
@@
A completion flow for $G$ is a $G$-flow $X$ containing a completion point, some $x\in X$ such that for any $\eta\in\widehat{G}$, the point $x\cdot\eta\in X$ has dense $G$-orbit, equivalently dense $\widehat{G}-$orbit. One of the main theorems from [Zuc19] is that if $G=\mathrm{Aut}(\mathbf{K})$ for $\mathbf{K}$ a Fraïssé structure which admits a big Ramsey structure, then $G$ admits a completion flow $X$ which is universal, i.e. which admits a $G$-map onto any other completion flow $Y$. For $G$ of this form, we further have that the universal completion flow is both metrizable and unique up to isomorphism.
@@
Let $f\in\mathrm{Emb}((\Sigma^{*},\preceq),(\Sigma^{*},\preceq))$. Let $g\in\mathrm{Shape}\allowbreak(\Sigma^{*},\allowbreak\Sigma^{*})$ and sequence $(N_{i})_{i\in\omega}$ be given by Lemma 8.3. Put $f^{\prime}=f\circ g\circ d$. Then for every poset-diary $S$ and every $\ell<\sup_{u\in S}|u|$, we have that $(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)\cong(\overline{f^{\prime}[S]}_{N_{3\ell}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$.
@@
$(\overline{S}_{0},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)=(\overline{f^{\prime}[S]}_{N_{0}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq).$
Now assume that $(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ is isomorphic to $(\overline{f^{\prime}[S]}_{N_{3\ell}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$. We define the function $\mu$ on $\overline{S}_{\ell+1}$ via
@@
We claim that $\mu\colon(\overline{S}_{\ell+1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)\to(\overline{f^{\prime}[S]}_{N_{3\ell+3}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ is an isomorphism:
@@
Fix $f\in\mathrm{Emb}(\mathbf{P},\mathbf{P})$. Let $\psi\colon\mathbf{P}\to{(\Sigma^{*},\preceq)}$ be obtained by Theorem 5.2. Let $\eta\colon{(\Sigma^{*},\preceq)}\to\mathbf{P}$ be an embedding (which exists since $\mathbf{P}$ is universal). Now $\psi\circ f\circ\eta$ is an embedding ${(\Sigma^{*},\preceq)}\to{(\Sigma^{*},\preceq)}$. Let $g\colon{(\Sigma^{*},\preceq)}\to{(\Sigma^{*},\preceq)}$ and $(N_{i})_{i\in\omega}$ be obtained by the application of Lemma 8.3 on $\psi\circ f\circ\eta$. Put $f^{\prime}=\psi\circ f\circ\eta\circ g\circ d$. We claim that for every poset-diary $S$, we have that $\tau(f^{\prime}[S])=S$. By Lemma 8.4, we know that for every $\ell<\sup_{u\in S}|u|$, we have that $(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)\cong(\overline{f^{\prime}[S]}_{N_{3\ell}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$. By Proposition 4.1, for every $0<\ell<\sup_{u\in S}|u|$, there is exactly one difference between $(\overline{S}_{\ell-1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ and $(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$. Consequently there is only one interesting level $\ell^{\prime}$ of $\overline{f^{\prime}[S]}$ between $N_{3\ell}$ and $N_{3\ell+3}$, and $(\overline{f^{\prime}[S]}_{N_{\ell^{\prime}}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)\cong(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$ while $(\overline{f^{\prime}[S]}_{N_{\ell^{\prime}+1}},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)\cong(\overline{S}_{\ell+1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\penalty 10000,\allowbreak\preceq\penalty 10000,\allowbreak\trianglelefteq)$. We further note that by the construction of $g$, the map $\tau_{f^{\prime}[S]}\circ f^{\prime}$ must be the identity on $S$. We can thus put $\varphi=\eta\circ g\circ d$. ∎
@@
Given a finite poset $\mathbf{A}$, the fact that the big Ramsey degree of $\mathbf{A}$ is exactly $|T^{\mathrm{lab}}(\mathbf{A})|=|T(\mathbf{A})|\cdot|\mathrm{Aut}(\mathbf{A})|$ follows from Corollary 7.8 and Theorem 8.1. To conclude that $\mathbf{P}$ admits a big Ramsey structure, we observe that the colourings $\chi_{\mathbf{A}}$ as $\mathbf{A}$ ranges over all finite posets satisfy the hypotheses of Theorem 7.1 from [Zuc19]. Theorem 1.6 from [Zuc19] then shows that $\mathrm{Aut}(\mathbf{P})$ admits a metrizable universal completion flow. ∎
@@ 9.1. Comparison to big Ramsey degrees of the order of rationals
Definition 9.1 (Devlin embedding type [Dev79], see also [Tod10], Definition 6.9).
A set $S\subseteq\{\mathrm{L},\mathrm{R}\}^{*}$ is called a Devlin embedding type, if no member of $S$ extends any other and precisely one of the following two conditions is satisfied for every level $\ell$ with $0\leq\ell<\sup_{w\in S}|w|$:
@@
    $\begin{split}\qquad\overline{S}_{\ell+1}&=\{z\in\overline{S}_{\ell}:z\mathbin{<_{\mathrm{lex}}}w\}^{\frown}\mathrm{L}\\ &\qquad\cup\{w^{\frown}\mathrm{L},w^{\frown}\mathrm{R}\}\\ &\qquad\cup\{z\in\overline{S}_{\ell}:w\mathbin{<_{\mathrm{lex}}}z\}^{\frown}\mathrm{L}.\end{split}$
@@
Theorem 9.3 (Devlin [Dev79], see also [Tod10]).
@@
Any infinite Devlin tree whose leaves code the rational order is a big Ramsey structure for the rationals [Zuc19].
@@ 9.2. The triangle-free graph
We outline how the techniques introduced in this paper can yield a particularly compact characterisation of the big Ramsey degrees of the generic triangle-free graph. This is a special case of the main result of [BCD+24] (see Example 6.2.10). However, we now give a more compact description which is similar to Definition 1.5.
We fix an alphabet $\Sigma=\{0,1\}$ and denote by $\Sigma^{*}$ the set of all finite words in the alphabet $\Sigma$, and by $|w|$ the length of the word $w$ (whose letters are indexed by natural numbers starting at 0). We will denote the empty word by $\emptyset$.
We consider the following triangle-free graph $\mathbf{G}^{\triangle}$ introduced in [Hub25].
@@
    $\overline{S}_{i+1}=\{z\in\overline{S}_{i}\setminus\{w\}:z\perp w\}^{\frown}0\cup\{z\in\overline{S}_{i}\setminus\{w\}:z\not\perp w\}^{\frown}1.$
@@
    $\overline{S}_{i+1}=(\overline{S}_{i}\setminus\{0^{i}\})^{\frown}0\cup\{{0^{i}}^{\frown}1\}.$
@@
Tringle-free diares use 0 to represent non-edges and $1$ to represent edges using the “passing number” representation which has been known for the Rado graph [Tod10, LSV06]. It is again interesting to understand triangle-free diaries in the context of their level structures. While for the Rado graph the level structure is trivial and holds no information except for the lexicographic order, in the triangle-free graph it distinguishes two types of vertices. If vertex has no neighbour defined yet (it is a word $0\cdots 0$), it may split into vertices connected by an edge. There is at most one such vertex in a level strcture. On the other hand, if vertex has a neighbour already defined, splitting it to two vertices connected by an edge would close a triangle. Consequently, it is an interesting event when the first neighbour of a vertex is introduced. Moreover, a given pair of vertices may or may not have a common neighbour. If common neighbour exists, there will be no edges between their respective successors. This is recorded by relation $\perp$ and the corresponding interesting event.
Given a triangle-free graph $\mathbf{H}$, we let $T^{\triangle}(\mathbf{H})$ be the set of all triangle-free-types $S$ such that the structure induced by $\mathbf{G}^{\triangle}$ on $L(S)$ is isomorphic to $\mathbf{H}$. We can now recover the following result of [BCD+24].
Theorem 9.8 ([BCD+24]).
@@ 10. Concluding remarks
The techniques presented in this paper are quite general and can be applied to a relatively wide class of structures in finite relational languages with relations of arity at most 2, defined by special triangle conditions isolated in [BCH+21]. Important examples of such structures are metric spaces with finitely many distances, or metric spaces associated to primitive metrically homogeneous graphs of finite diameter from Cherlin’s catalogue of metrically homogeneous graphs [Che22, ABWH+26].
However, forbidding a configuration on 4 or more vertices (such as in the universal homogeneous $K_{4}$-free graph) will run into a limitation of regular trees used here. This can be overcome in two ways. Either a Ramsey theorem for coding trees can be proved [Dob20a, Dob23b, Zuc22, BCD+24] or a theorem for trees with finite but unbounded branching can be used [BCD+23] to represent a more regular tree constructed from a kind of amalgamation of all coding trees corresponding to all possible enumerations of the structure, as used in this paper.
Structures with relations of arity 3 and more require trees with unbounded branching together with more involved trees of weak types. Upper bounds for structures with very simple restrictions can be obtained using the product form of the Milliken tree theorem [BCH+22, BCdR+24], while more general results can be obtained using trees with a successor operation [BCD+23], see also the announcement in [ABC+23]. Even the characterisation of big Ramsey degrees of 3-uniform hypergraphs shows new and somewhat surprising phenomena, and is a topic of a follow-up paper.
The actual counts of diaries defined in this paper, as well as for diares of $K_{4}$-free graphs, are computed in [Vod25, HKVZ25].